OpenStudy (anonymous):

I really need help Will fan and medal A scale on a map shows that 2 inches = 25 miles. Part A How many inches on the map represent 60 miles? Enter your answer in the box. inches Part B How many miles are represented by 1 4 inch on the map? Enter your answer in the box. miles

OpenStudy (bibby):

set up proportions, 2:25=x:60

OpenStudy (anonymous):

ok @bibby

OpenStudy (anonymous):

now what? @bibby

OpenStudy (bibby):

solve for the unknown

OpenStudy (anonymous):

ya but how? @bibby

OpenStudy (bibby):

cross multiplying

OpenStudy (anonymous):

ok so 2 x 60 = 120. what do I do with x and 25? @bibby

OpenStudy (bibby):

\(\dfrac{2in}{25mi}=\dfrac{xin}{60mi}\)

OpenStudy (bibby):

so multiply the 25 and x

OpenStudy (anonymous):

how? @bibby

OpenStudy (bibby):

25*x=25x...

OpenStudy (bibby):

so we have 25x=60*2

OpenStudy (anonymous):

so it's 25x = 120 then you divide both sides by 25 right? @bibby

OpenStudy (bibby):

aye aye captain

OpenStudy (bibby):

btw stop tagging me

OpenStudy (anonymous):

x = 4.8 right?

OpenStudy (bibby):

mhm

OpenStudy (anonymous):

Great! thanks!

OpenStudy (anonymous):

think you can help me with part B?

OpenStudy (bibby):

yep same thing, set up a proportion

OpenStudy (bibby):

\[\frac{ ?\in }{ ?mi }=\frac{ ?\in }{ ?mi }\]

OpenStudy (anonymous):

ok so 2:25 = x = 1/4 right?

OpenStudy (bibby):

not exactly. not sure what you're doing on the righty hand side

OpenStudy (anonymous):

oh ok

OpenStudy (bibby):

remember its inches /miles

OpenStudy (anonymous):

oh ya

OpenStudy (anonymous):

ok what now

OpenStudy (bibby):

need another clue?

OpenStudy (bibby):

show me your proportion

OpenStudy (anonymous):

\[\frac{ 1 }{ 4 } = \frac{ 2 }{ 25 }\]

OpenStudy (bibby):

the 1/4 os a fraction, it's a single unit of inches. 2 inches for every 25 miles 1/4 inch for ???? mile

OpenStudy (bibby):

it might help to rewrite it in decimal

OpenStudy (bibby):